# 11footnotetext: {*} denotes Equal contribution

### v1

@@ Abstract
Large Language Models (LLMs) have dramatically advanced AI applications, yet their deployment remains challenging due to their immense inference costs. Recent studies ameliorate the computational costs of LLMs by increasing their activation sparsity but suffer from significant performance degradation on downstream tasks. In this work, we introduce a new framework for sparsifying the activations of base LLMs and reducing inference costs, dubbed Contextually Aware Thresholding for Sparsity (CATS). CATS is relatively simple, easy to implement, and highly effective. At the heart of our framework is a new non-linear activation function. We demonstrate that CATS can be applied to various base models, including Mistral-7B and Llama2-7B, and outperforms existing sparsification techniques in downstream task performance. More precisely, CATS-based models often achieve downstream task performance within 1-2% of their base models without any fine-tuning and even at activation sparsity levels of 50%. Furthermore, CATS-based models converge faster and display better task performance than competing techniques when fine-tuning is applied. Finally, we develop a custom GPU kernel for efficient implementation of CATS that translates the activation of sparsity of CATS to real wall-clock time speedups. Our custom kernel implementation of CATS results in a $\sim$15improvement in wall-clock inference latency of token generation on both Llama-7B and Mistral-7B.
@@ 1 Introduction
LLMs have demonstrated remarkable success across a variety of fields (Devlin et al., 2018; Brown et al., 2020; Achiam et al., 2023; Brohan et al., 2023). However, the scientific progress achieved by these models comes with significant costs. The training of GPT-3 is estimated to have consumed over 3,000,000 GPU-hours and emitted three thousand times the CO2 equivalent of a round-trip flight from San Francisco to New York (Patterson et al., 2021). Furthermore, inference costs often eclipse training costs for models that serve trillions of queries. As such, there is significant interest in reducing the inference costs of LLMs while preserving task performance.
@@
MoE techniques activate only a subset of parameters at each inference stage, thereby reducing memory and computational requirements compared to using the entire model. Prevailing implementations of MoE techniques introduce many multi-layer perceptrons (MLPs; the “experts”) and dynamically select which experts to multiply with the hidden vector. This selection is performed by a “router”–a small neural network trained to determine the appropriate experts to activate based on the input (Lewis et al., 2021; Rajbhandari et al., 2020).
Concurrently, recent work has observed that activations in the MLP blocks of LLMs are sparse Liu et al. (2023b); Mirzadeh et al. (2023). This implies that only a few rows (or columns) of the corresponding weight matrices are required for the forward pass. Intuitively, if we could predict a priori which elements of the weight matrices were unnecessary via an oracle, we could obviate their respective computations. This is thematically similar to MoE approaches: the activated neurons of the weight matrices can be viewed as activated “experts” and the oracle can be seen as the “router.”
@@
Figure 1: Histogram of post-MLP activations different layers in different models. Subfigures (a), (b), and (c) correspond to Layers 0, 15, and 31 in Llama2-7B, respectively. Subfigures (d), (e), and (f) correspond to Layers 0, 15, and 31 in Mistral-7B, respectively. The absolute threshold indicates 50% sparsity, where values smaller than the threshold are considered negligible in our technique and thus zeroed out.
@@
    We motivate the development of a new sparsification procedure based on a novel activation function, dubbed CATS (for Contextually Aware Thresholding for Sparsity), motivated by an empirical evaluation of activation distributions (Figure 1). Crucially, CATS allows for a controllable level of sparsity.
@@
    We demonstrate that, without any fine-tuning, CATS can be applied to various base models, including Mistral-7B and Llama2-7B, and achieves comparable downstream task performance even at sparsity levels as high as 50%.
@@
    We demonstrate that, with fine-tuning, CATS outperforms existing state-of-the-art sparsification techniques in downstream task performance at the same sparsity level and number of fine-tuning steps.
@@
    We provide a custom GPU kernel implementation that exploits the sparsity of CATS and achieves a $\sim$15improvement in wall-clock inference latency of token generation over the dense base models.
@@ 3 Background
Motivation: As described in Section 1, MoE models selectively activate expert subnetworks via a trained router. Crucially, we may view the rows (or columns) of MLP layers as experts in an MoE model. To identify the layers most likely to benefit from this MoE view (where many activations can be zeroed), we examine the activations of different layers in LLMs. Figure 1 demonstrates that activations of the Gated-MLP layers tend to concentrate around zero across different LLMs. This behavior suggests that many neurons of MLP layers minimally affect the output in future operations.
Gated-MLP Blocks: We now describe the components of LLMs that our work aims to accelerate: the Gated-MLP blocks. Gated-MLP blocks are commonly used in LLMs, including in the Llama2 family of models, Mistral-7B, and Gemma. A Gated-MLP block consists of several fully-connected layers and performs the following computation:
@@
where $x\in{\mathbb{R}}^{b\times d}$, $W_{\text{up}}\in{\mathbb{R}}^{m\times d}$, $W_{\text{gate}},W_{\text{up}}\in{\mathbb{R}}^{d\times m}$, * indicates elementwise multiplication, and
@@
Crucially, the operation $\text{SiLU}(xW_{\text{gate}})$ can be viewed as the router in an MoE model. Under this lens, the columns of $W_{\text{up}}$ and the rows of $W_{\text{down}}$ are the experts. If $\text{SiLU}(x)$ is always binary, i.e., $1$ or 00, it would turn on/off elements of the remaining computation (multiplication by $W_{\text{up}}W_{\text{down}}$). When $\text{SiLU}(x)$ is not binary, it can be viewed as a “soft” router that weighs the experts by different amounts.
@@ 4 Method: Contextually-Aware Thresholding for Sparsification (CATS)
We now describe CATS, a framework to accelerate the Gated-MLP blocks of LLMs. The CATS framework proposes a new activation function and exploits the sparsity induced by this activation. In Section 5, we apply CATS to Mistral-7B and Llama2-7B and show that CATS-based models still exhibit significant activation sparsity, even when fine-tuned.
@@ 4.1 Stage 1: Determining Cutoff Threshold
We assume we are given a desired sparsity level $k$ (e.g., 70%) as input. For each Gated-MLP block in the LLM, we compute the activations over a random subset of the training data. We then compute the cutoff threshold as the $k$th percentile of the resulting values.
More formally, the cutoff threshold $t$ is
@@
where $F$ is the empirical CDF of activations’ absolute values for the given MLP block.
Figure 1 shows histograms of the absolute values of activations of the different MLP block in different models over the RefinedWeb dataset (Penedo et al., 2023). A sparsity level of 70% corresponds to a threshold of approximately 0.15; different sparsity levels correspond to different thresholds. We note that these thresholds are chosen and fixed before any further fine-tuning.
@@ 4.2 Stage 2: Sparsifying Gate-MLP Blocks
$\text{CATS}_{t}(\mathbf{x})_{j}:=\begin{cases}x_{j},&\text{if }|x_{j}|\geq t\\ 0,&\text{if }|x_{j}|<t\end{cases}$ (4)
Here, $t$ is the sparsification threshold and $x_{j}$ is the $j$-th element of vector $\mathbf{x}$, respectively.
@@
Intuitively, the resulting model zeros out activations which were likely to be close to 00 because their corresponding inputs were small. This procedure results in a trained model whose activations are sparse and whose performance may then be evaluated. We empirically validate that this procedure results in a model with sparsity level approximate $k$, even after fine-tuning, in Appendix C.
@@ 4.3 Custom Kernel Design
The previous subsections describe the procedure for sparsifying LLM’s activations, obviating computations, and reducing the required number of floating point operations (FLOPs) in each MLP block. We now translate the reduction in FLOPs to a reduction in actual wall-clock latency and increase in generation throughput via a custom GPU kernel.
@@
We focus on reducing the latency of the MLP blocks by reducing memory accesses because the MLP blocks are known to be memory-bound during inference (Kim et al., 2023). As shown in Line 5 of the Custom GPU Kernel 1, we first fuse the element-wise multiplication of $v[\texttt{Mask}]$ into each tiling of $xW_{\text{up}}[\texttt{Mask}]$ where $v$ is the hidden vector after the SiLU activations and Mask is a binary mask that labels the elements of $v$ with large absolute value. This fusion saves memory operations that would be necessary for storing and loading $x_{1}$ several times. We then directly use Mask to control which parts of the weight matrices $W_{\text{up}}$ and $W_{\text{down}}$ to load, instead of using the compressed indices directly as in Zhang et al. (2023) This further improves the kernel speed because it avoids expensive synchronization operations. We demonstrate the success of this custom GPU kernel at reducing the inference latency of CATS-based models as the sparsity increases in Section 5.2.
@@ 5 Experiments
In this section, we describe the experiments with which we assess the performance of CATS. We first describe the experimental details that are common to all experimental settings. We then describe experiments on downstream task performance. Finally, we measure CATS’ effect on wall-clock time inference when implemented with the custom GPU kernel from Section 4. We find that CATS-based models outperform the baseline models and their ReLUfication versions in downstream task performance, with or without fine-tuning, and can exploit their sparsity for wall-clock inference time speedups over the base models.
@@
CATS-based Models: For a given base model, we train three CATS-based variants that exhibit different sparsity levels in the MLP blocks: 50%, 70%, and 90% activation sparsity. We call these models CATS 50%, CATS 70%, and CATS 90%, respectively, where the base models are clear from context.
@@
Datasets: For the downstream task performance experiments, we use the OpenBookQA, ARC_Easy, Winogrande, HellaSwag, ARC_Challenge, PIQA, BoolQ, and SCI-Q datasets from the Eleuther AI Evaluation Harness (Gao et al., 2023) as in Mirzadeh et al. (2023) for ease of comparison; these tasks were originally chosen to measure various abilities of the models across various domains, such as reading comprehension and reasoning. For the latency experiments, we assess the wall-clock inference time on the RefinedWeb test dataset (Penedo et al., 2023).
Computational Environment: All experiments were run on a single machine with 8 L40S GPUs. Latency experiments were run on a single L40S GPU as each 7B base model was able to fit in a single GPU RAM when performing inference in brain float 16 (BF16) or floating point 16/32 (FP16/32) precision. We used DeepSpeed with BF16 precision to manage the high memory overhead during training. We also employed LoRA and targeted 1% of the parameters (Query and Key in attention modules, $W_{\text{gate}}$, and $W_{\text{down}}$) in the fine-tuning experiments. During inference, we used the transformers v4.36.2 HuggingFace library, PyTorch v2.1.2, and CUDA v12.1. We used Triton v2.1.0 for our GPU kernels. All experiments were run in FP32 precision; changing this to FP16 did not materially affect results. All of our code, including a one-line script to set up an environment and reproduce all of our results, is available in the supplementary material.
@@ 5.1 Downstream Task Performance
Table 1: Downstream task performance of base models, CATS-based models for varying levels of sparsity, and ReLUfication across different benchmark datasets. The CATS versions of the base models demonstrate comparable performance to original models at high sparsity levels, e.g., 50%. At higher sparsity levels, CATS still outperforms ReLUfication.
CATS-based models perform comparably to the base models and outperform ReLUfication in zero-shot accuracy with “general” fine-tuning: In this setting, CATS is applied the base models Llama-7B and Mistral-7B. All models are then fine-tuned on the RefinedWeb dataset Penedo et al. (2023); their downstream performance is then measured on the 8 evaluation datasets. We emphasize that the dataset upon which the models are fine-tuned is different from the evaluation datasets in this setting. Figure 2 demonstrates our results. We note several key observations:
@@
    CATS-50% demonstrates performance comparable to the base models, even without any fine-tuning (0 fine-tuning steps). This is in contrast with ReLUficiation, which demonstrates poor performance without fine-tuning.
@@
Figure 2: Downstream task performance of the base model, CATS models with different sparsity levels, and ReLUfication versus number of fine-tuning steps on the RefinedWeb dataset applied to Mistral-7B (left) and Llama2-7B (right). The CATS models demonstrate faster convergence and better fine-tuning efficiency than the ReLUfication variants. Furthermore, CATS-50% and CATS-70% demonstrate comparable performance to the base models without any fine-tuning (0 fine-tuning steps).
CATS-based models perform comparably to the base models and outperform ReLUfication in zero-shot accuracy with task-specific fine-tuning: In this setting, the CATS prescription is applied to Mistral-7B. All variants are then fine-tuned for 10 epochs on the training data and evaluated on test dataset for the Cola, SST2, and BoolQ datasets. Table 2 demonstrates our results. Our observations are similar to those for “general” fine-tuning:

### v4

# CATS: Contextually-Aware Thresholding for Sparsity in Large Language Models

@@ Abstract
The dramatic improvements in Large Language Models (LLMs) come at the cost of increased computational resources for inference. Recent studies ameliorate the computational costs of LLMs by increasing their activation sparsity but suffer from significant performance degradation on downstream tasks. In this work, we introduce a new framework for sparsifying the activations of LLMs and reducing inference costs, dubbed Contextually Aware Thresholding for Sparsity (CATS). CATS is a relatively simple algorithm that is easy to implement and highly effective. At the heart of our framework is a new non-linear activation function. We demonstrate that CATS can be applied to various models, including Mistral-7B and Llama2-7B &\&& 13B, and outperforms existing sparsification techniques across multiple tasks. More precisely, CATS-based models achieve downstream task performance within $\sim$99of their base models at 50% activation sparsity, even without fine-tuning. Moreover, with fine-tuning that targets only 1% of the parameters, CATS-based models not only converge faster but also achieve better task performance than competing techniques. Finally, we develop a custom GPU kernel for efficient implementation of CATS that translates the activation sparsity of CATS to real wall-clock time speedups. Our custom kernel implementation of CATS results in a $\sim$15improvement in wall-clock inference latency of token generation. We release our code, experiments, and datasets at https://github.com/ScalingIntelligence/CATS.
11footnotetext: These authors contributed equally as co-first authors.
@@ 1 Introduction
LLMs have demonstrated remarkable success across a variety of fields (Devlin et al., 2018; Brown et al., 2020; Achiam et al., 2023; Brohan et al., 2023), however, this progress comes with significant computational costs. The training of GPT-3 is estimated to have consumed over 3,000,000 GPU-hours and emitted three thousand times the CO2 equivalent of a round-trip flight from San Francisco to New York (Patterson et al., 2021). Furthermore, inference costs often eclipse training costs for models that serve trillions of queries. As such, there is significant interest in reducing the inference costs of LLMs while preserving task performance.
@@
MoE techniques activate only a subset of parameters at each inference stage, thereby reducing memory and computational requirements compared to using the entire model. Prevailing implementations of MoE techniques introduce many multi-layer perceptrons (MLPs; the “experts”) and dynamically select which experts to multiply with the hidden vector. This selection is performed by a “router”—a small neural network trained to determine the appropriate experts to activate based on the input (Lewis et al., 2021; Rajbhandari et al., 2020).
Concurrently, recent work has observed that activations in the MLP blocks of LLMs are sparse (Liu et al., 2023b; Mirzadeh et al., 2023). This implies that only a few rows (or columns) of the corresponding weight matrices are required for the forward pass. Intuitively, if we could predict a priori which elements of the weight matrices were unnecessary via an oracle, we could obviate their respective computations. This is thematically similar to MoE approaches: the activated neurons of the weight matrices can be viewed as activated “experts” and the oracle can be seen as the “router.”
@@
Figure 1: Histograms of post-MLP activations of different layers in different models. Subfigures (a), (b), and (c) correspond to Layers 0, 15, and 31 in Llama2-7B, respectively. Subfigures (d), (e), and (f) correspond to Layers 0, 15, and 31 in Mistral 7B, respectively. The absolute threshold indicates 50% sparsity, where values smaller than the threshold are considered negligible in our technique and thus zeroed out.
@@
    We introduce a new sparsification procedure based on a novel activation function, dubbed CATS (for Contextually Aware Thresholding for Sparsity), motivated by an empirical evaluation of activation distributions (Figure 1). Crucially, CATS allows for a controllable level of sparsity.
@@
    We demonstrate that, without any fine-tuning, CATS can be applied to various models, including Mistral-7B and Llama2-7B &\&& 13B, and achieves comparable downstream task performance even at sparsity levels as high as 50%.
@@
    We demonstrate that, with parameter-efficient fine-tuning, CATS outperforms an existing sparsification technique in downstream task performance at the same sparsity level and number of fine-tuning steps.
@@
    We provide a custom GPU kernel implementation that exploits the sparsity of CATS and achieves a $\sim$15improvement in wall-clock inference latency of token generation over the dense models.
@@ 3 Background
Motivation: As described in Section 1, MoE models selectively activate expert subnetworks via a trained router. Crucially, we may view the rows (or columns) of MLP matrices as experts in an MoE model. To identify the layers most likely to benefit from this MoE perspective (where many activations can be zeroed), we examine the activations of different layers in LLMs. Figure 1 demonstrates that activations of the Gated-MLP layers tend to concentrate around zero across different LLMs. This behavior suggests that many neurons of MLP layers minimally affect the output in future operations.
Gated-MLP Blocks: We now describe the components of LLMs that our work aims to accelerate: the Gated-MLP blocks. They are commonly used in LLMs, including in the Llama2 family of models, Mistral-7B, and Gemma. A Gated-MLP block consists of several fully-connected layers and performs the following computation:
@@
where $x\in{\mathbb{R}}^{b\times d}$, $W_{\text{gate}},W_{\text{up}}\in{\mathbb{R}}^{d\times m}$, $W_{\text{down}}\in{\mathbb{R}}^{m\times d}$, * indicates element-wise multiplication, and
@@
Crucially, the operation $\text{SiLU}(xW_{\text{gate}})$ can be viewed as the router in an MoE model. Through this lens, the columns of $W_{\text{up}}$ and the rows of $W_{\text{down}}$ are the experts. If $\text{SiLU}(x)$ is always binary, i.e., $1$ or 00, it would turn on/off elements of the remaining computation (multiplication by $W_{\text{up}}W_{\text{down}}$). When $\text{SiLU}(x)$ is not binary, it can be viewed as a “soft” router that weighs the experts by different amounts.
@@ 4 Method: Contextually-Aware Thresholding for Sparsification (CATS)
We now describe CATS, a framework to accelerate the Gated-MLP blocks of LLMs. The CATS framework proposes a new, simple activation function and exploits the sparsity induced by this activation. In Section 5, we apply CATS to Mistral-7B and Llama2-7B and show that CATS-based models still exhibit significant activation sparsity, even when fine-tuned.
@@ 4.1 Stage 1: Determining Cutoff Threshold
We assume we are given a desired sparsity level $k$ (e.g., 70%) as input. For each Gated-MLP block in the LLM, we compute the activations over a random subset of the training data, limited to only 500 data points. We then compute the cutoff threshold as the $k$th percentile of the resulting values.
Concretely, the cutoff threshold $t$ is
@@
where $F$ represents the empirical cumulative distribution function of the absolute values of the activations for the given MLP block.
Figure 1 shows histograms of the absolute values of activations for different MLP blocks in various models on the RefinedWeb dataset (Penedo et al., 2023). A sparsity level of 70% corresponds to a threshold of approximately 0.15; different sparsity levels correspond to different thresholds. We note that these thresholds are chosen and fixed before any further fine-tuning.
@@ 4.2 Stage 2: Sparsifying Gate-MLP Blocks
$\text{CATS}_{t}(\mathbf{x_{j}}):=\begin{cases}x_{j},&\text{if }|x_{j}|\geq t\\ 0,&\text{if }|x_{j}|<t\end{cases}$ (4)
Here, $t$ is the sparsification threshold and $x_{j}$ is the $j$-th element of the vector $\mathbf{x}$, respectively.
@@
Intuitively, the resulting model zeros out activations that are likely to be close to 00 because their corresponding inputs were small. This procedure produces a trained model with sparse activations, whose performance can then be evaluated. We empirically validate that this procedure results in a model with an approximate sparsity level of $k$, even after fine-tuning, as detailed in Appendix C.
@@ 4.3 Custom Kernel Design
The previous subsections describe the procedure for sparsifying LLM’s activations, obviating unnecessary computations, and reducing the required number of floating point operations (FLOPs) in each MLP block. We now translate this reduction in FLOPs to a reduction in actual wall-clock latency and increase in generation throughput via a custom GPU kernel.
@@
We focus on reducing the latency of the MLP blocks by reducing memory accesses, as the MLP blocks are known to be memory-bound during inference when the batch size is small (Kim et al., 2023). As shown in Line 4 of the Custom GPU Kernel 1, we first fuse the element-wise multiplication of $v[\texttt{Mask}]$ into each tiling of $xW_{\text{up}}[\texttt{Mask}]$. Here, $v$ represents the hidden vector after applying the SiLU activation function, and Mask denotes a binary mask identifying elements of $v$ with large absolute values. This fusion saves memory operations that would be necessary for storing and loading $x_{1}$ several times. We then directly use Mask to control which parts of the weight matrices $W_{\text{up}}$ and $W_{\text{down}}$ to load, instead of using the compressed indices directly as in Zhang et al. (2023) This further improves the kernel speed because it avoids expensive synchronization operations. In Section 5.3, we demonstrate how our custom GPU kernel effectively reduces the inference latency of CATS-based models as sparsity increases.
@@ 5 Experiments
In this section, we describe the experiments with which we assess the performance of CATS. We first describe the experimental details that are common to all experimental settings. We then describe experiments on downstream task performance. Finally, we measure CATS’ effect on wall-clock time inference when implemented with the custom GPU kernel from Section 4. We find that CATS-based models outperform their ReLUfication versions in downstream task performance, with or without fine-tuning, and can exploit their sparsity for wall-clock inference time speedups over the base models.
@@
CATS-based Models: For a given base model, we train three CATS-based variants that exhibit different sparsity levels in the MLP blocks: 50%, 70%, and 90% activation sparsity. We call these models CATS-50%, CATS-70%, and CATS-90%, respectively, where the base models are clear from context.
@@
Datasets: For the downstream task performance experiments, we use the OpenBookQA, ARC_Easy, Winogrande, HellaSwag, ARC_Challenge, PIQA, BoolQ, and SCI-Q datasets from the Eleuther AI Evaluation Harness (Gao et al., 2023) as in Mirzadeh et al. (2023) for ease of comparison; these tasks were originally chosen to measure various abilities of the models across various domains, such as reading comprehension and reasoning. For the latency and generation task experiments, we assess the wall-clock inference time and perplexity score on the RefinedWeb test dataset (Penedo et al., 2023).
Computational Environment: All experiments were run on a single machine with 8 L40S GPUs. Latency experiments were run on a single L40S GPU as each 7B base model was able to fit in a single GPU RAM when performing inference in brain float 16 (BF16) or floating point 16/32 (FP16/32) precision. We used DeepSpeed (Rasley et al., 2020) with BF16 precision to manage the high memory overhead during training. We also employed Low-Rank Adaptation (LoRA) (Hu et al., 2021) and targeted 1% of the parameters (Query and Key in attention modules, $W_{\text{gate}}$, and $W_{\text{down}}$) in the fine-tuning experiments. During inference, we used the transformers v4.36.2 HuggingFace library, PyTorch v2.1.2, and CUDA v12.1. We used Triton v2.1.0 for our GPU kernels. All experiments were run in FP32 precision; changing this to FP16 did not materially affect results. All of our code, including a one-line script to set up an environment and reproduce all of our results, is available in the supplementary material.
@@ 5.1 Downstream Task Performance
CATS-based models perform more comparably to the base model and increasingly outperform ReLUfication as the model size increases in zero-shot accuracy: Similar to the previous evaluation, we assess the performance of various CATS configurations and a baseline Llama-13B model across eight downstream tasks without any fine-tuning. Notably, the performance degradation of CATS-50% decreased from 1.46% at 7B to 0.65% at 13B. Conversely, the performance gap between CATS-50% and ReLUfication increased from 29.08% at 7B to 34.55% at 13B. We present the detailed results in Appendix E.
@@
Table 1: Downstream task performance of base models, CATS-based models at varying levels of sparsity, and ReLUfication across different benchmark datasets. The CATS versions of the base models demonstrate comparable performance to original models at high sparsity levels, e.g., 50%. At higher sparsity levels, CATS still outperforms ReLUfication.
CATS-based models perform comparably to the base models and outperform ReLUfication in zero-shot accuracy with “general” fine-tuning: In this setting, CATS is applied the base models Llama-7B and Mistral-7B. All models are then fine-tuned using LoRA (Hu et al., 2021), targeting only 1% of the parameters, on the RefinedWeb dataset Penedo et al. (2023). Their downstream performance is subsequently measured across 8 evaluation datasets. We emphasize that the dataset upon which the models are fine-tuned is different from the evaluation datasets in this setting. Figure 2 demonstrates our results. We note several key observations:
@@
    CATS-50% demonstrates performance comparable to the base models, even without any fine-tuning. This is in contrast with ReLUficiation, which demonstrates poor performance without fine-tuning.
@@
Figure 2: Downstream task performance of the base model, CATS models with different sparsity levels, and ReLUfication across varying numbers of fine-tuning steps on the RefinedWeb dataset applied to Mistral-7B (left) and Llama2-7B (right). The CATS models exhibit faster convergence and greater fine-tuning efficiency than the ReLUfication variants. Furthermore, CATS-50% and CATS-70% demonstrate comparable performance to the base models without any fine-tuning (0 fine-tuning steps).
CATS-based models perform comparably to the base models and outperform ReLUfication with task-specific fine-tuning but without “general“ fine-tuning: In this setting, the CATS prescription is applied to Mistral-7B. All variants are then fine-tuned for 10 epochs on the training data and evaluated on test dataset for the Cola, SST2, and BoolQ datasets. Table 2 demonstrates our results. Our observations are similar to those for “general” fine-tuning:
@@
5.2 Generation task performance
We now evaluate the generation task performance of the CATS-based models using the RefinedWeb dataset. Without any fine-tuning, CATS-50% exhibits only a slight increase in perplexity scores. The difference in perplexity scores between the base models and CATS-50% further diminishes after 1300 fine-tuning steps.
Specifically, for Llama-7B, the perplexity increases by just 1.06% when comparing the base model to CATS-50% without fine-tuning, and this difference narrows to 0.60% with minimal fine-tuning. Similarly, for Mistral-7B, the initial perplexity increase of 1.21% is reduced to a mere 0.45% after fine-tuning.
Model Base Model CATS 50% CATS 50% ReLUfication ReLUfication
Llama-7B 2.18 2.203 (+1.06%) 2.193 (+0.60%) 3.104 (+42.38%) 2.617 (+20.04%)
Mistral-7B 2.234 2.261 (+1.21%) 2.244 (+0.45%) 2.801 (+25.34%) 2.562 (+14.68%)
Table 3: Perplexity scores for different models and configurations. The best score is marked in bold and the second-best is underlined. Percentage differences relative to the base model are shown in parentheses. Models marked with are fine-tuned for 1300 steps.